OpenStudy (anonymous):

factor the polynomials completely. 18q^2 - 98n^2

OpenStudy (anonymous):

18q^2 - 98n^2 =2(9q^2-49n^2) GCF =2(3q-7n)(3q+7n) diff two squares

OpenStudy (saifoo.khan):

Right.

OpenStudy (anonymous):

ty so much again mandolino

OpenStudy (anonymous):

np :})